INAF queue observing, 2018 November 05-06

Observer: R. Carini, M. Faccini
Telescope Operator: S. Allenson
Support Astronomer: B. Rothberg
Instrument: LUCI1 (LS mode), LUCI2 stops working at 03:14 UT, we can not do MOS programs, MODS1

Summary

Night duration (12\xB0 el twilight) 1:15-12:45 UT = 11h 30m

Observing time: 10h 45min (94%)

Weather time loss: h m (0%)

technical time loss 0h 45m (6%)

Data Summary

At the sunset the sky is clear, wind about 10 m/s, humidity about 43%

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
dark Antoniucci Ves 691 + tell Luci1+luci2     luci1.0001-0015, luci2.0001-0015  
dark Polletta PLCK073 +tell luci1+luci2     luci1.0016-0025, luci2.0016-0025  
dark Nardini q2 J030449 +tell luci1+luci2     luci1.0026-0035, luci2.0026-0035  
Antoniucci telluric (Ves 691) HD 216532 luci1+luci2 ls 0.75, G200LoRes, zjSpec,HKspec   luci1.0039-0046, luci2.0046-0053  
Antoniucci Ves 691 luci11+luci2 ls 0.75, G200LoRes, zjSpec,HKspec 3m +3m luci1.0052-0059, luci2.0059-0066 completed
Antoniucci telluric (IRAS 02276+7225) HD 18326 luci1+luci2 ls 0.75, G200LoRes, zjSpec,HKspec   luci1.0063-0068, luci2.0070-0075  
Antoniucci IRAS 02276+7225 luci1+luci2 ls 0.75, G200LoRes, zjSpec,HKspec 6m +6m luci1.0073-0080, luci2.0080-0087 completed
Antoniucci telluric (V1180 Cas) HIP10024 luci1+luci2 ls 0.75, G200LoRes, zjSpec,HKspec   luci1.0084-0087  
Antoniucci V1180 Cas luci1 ls 0.75, G200LoRes, zjSpec,HKspec 0.7h luci1.0092-0099 completed
Antoniucci Agll354 luci1 ls 0.75, G200LoRes, zjSpec,HKspec 0.5h luci1.0104-111 completed
Antoniucci telluric ( Agll 354, D75b ) HIP 25410 luci1 ls 0.75, G200LoRes, zjSpec,HKspec   luci1.0117-0124  
Antoniucci D75b luci1 ls 0.75, G200LoRes, zjSpec,HKspec 0.7h luci1.0129-0136 completed
Nardini telluric (q5) HIP 13917 luci1 ls 1.0, G150Ks, Ks   luci1.0142-0143  
Nardini q5 SDSS J0303 luci1 ls 1.0, G150Ks, Ks 0.67h luci1.0148-0171 completed
Moretti A14 FKJ0355+00 mods1 red grating 3600 s mods1.0005-0007 completed (we take one more image due to variable seeing)
Antoniucci V900 mon mods1 ls 0.8", dual grating 20 m mods1r.0015-0018, mods1b.0002-0005 completed
Antoniucci V899 mon mods1 ls 0.8", dual grating 20 m mods1r.0022-0025, mods1b.0006-0009 completed
standard GD 71 mods1 dual grating; red grating   mods1r.0028-0032, mods1b.0010-0012; mods1r.0031-0032  
Calibration Moretti MODS slitflat (1,0", 5.0"), slitless flat, lamp (0.6", 1,0") red grating  
Calibrations Antoniucci MODS slitflat (0.8"), slitless flat, lamp (0.6") dual grating  
Calibration Antoniucci LUCI flat and arc    

Details

Calibrations: darks
  • Antoniucci tell Ves 691* luci1.20181106.0001-0005, luci2.20181106.0001-0005 (dit 2.75,ndit 3, lir, int)
  • Ves 691 target ==luci1.20181106.0006-0015, luci2.20181106.0006-0015==(dit 25 ndit 1, norm, mer; dit 15 ndit 1, norm, mer)
  • Polletta PLCK073 luci1.0016-0020, luci2.0016-0020 target (dit 120, ndit 1, mer, norm)
  • Polletta telluric luci1.0021-0025, luci2.0021-0025 (dit 5, ndit 6, lir, int) (it is ok also for tell HIP10024 for Antoniucci)
  • Nardini q2, J030449 luci1.0026-0030, luci2.0026-0030 target (dit 110, ndit 1,mer,norm)
  • Nardini telluric (dit 5, ndit 3, lir, int ) luci1.0031-0035, luci2.0031-0035

12:16 TO opens the dome; AO check

01:10 start with Antoniucci program, telluric for Ves 691
  • 01:12 check copointing
  • 01:18 send again the script
  • acq luci1.0036-0038,luci2.0043-0045
  • seeig about 1.8''
  • sci luci1.0039-0046, luci2.0046-0053

01:36 slew to the target Ves 691
  • 01:39 problems with the preset
  • 01:45 send again the script, now the preset is working, but there are some issues with the GCS of LUCI2
  • 02:03 send again the preset, the collimations works
  • acq luci1.0047-0051,luci2.0054-0058
  • sci luci1.0052-0059 ,luci2.0059-0066
  • 02:18 seeing about 1.2'' but instable

02:22 slew to telluric for IRAS02
  • acq luci1.0060-0062,luci2.00670069
  • 02:33 sci luci1.0063-0068,luci2.0070-0075
  • seeing about 1.7''

02:39 slew to IRas 02
  • acq luci1.0069-0072,luci2.0076-0079
  • seeing about 1.0''
  • sci luci1.0073-0080,luci2.0080-0087
  • seeing about 1.5''-1.6''

03:09 slew to telluric for V1180 Cas
  • 03:10 problem with the preset with LUCI2
  • 03:14 the TO send the telescope to the Zenith to try to resolve the isseu, but it doesn't work. LUCI2 is not available fot this night. We continue with LUCI1 only

03:24 send the script with luci1, but there is a problem with the GCS

03:27 send again the scipts, error, we send again and now it works
  • 03:35 seeing 2-2.2''
  • acq luci1.0081-0083
  • sci luci1.0084-0087

03:39 send the script for the target V1180 Cas
  • acq luci.0088-0091
  • 03:46 sci luci1.0092-0099
  • 04:10 seeing about 1.4''-1.6''
  • 04:18 seeing 1.2'' on the DIMM and 0.9'' on the FWHM
  • 04:30 seeing again about 1.5''

04:32 slew to Agll354
  • check pointing
  • 04:38 send again the script
  • 04:40 error again
  • 04:41 send again the script
  • acq luci1.0100-0103
  • seeing 1.6''
  • 04:49 sci luci1.104-
  • 04:56 seeing about 1.5''
  • seeing about 2.0''
  • seeing very variable between 2.00'' and 1.5''

05:39 slew to the telluric of Agll354, but it is too low, we change it
  • 05:41 check pointing
  • 05:49 acq HIP 25410 luci1.0113-0116
  • sci = luci1.0117-0124=

06:00 slew to D75b
  • acq luci1.0125-0128
  • sci luci.0129-0136

06:27 slew to tellurci HIP 13917 for Nardini's program
  • 06:29 check pointing
  • seeing variable between 1.2"-1.6"
  • acq luci1.0137-0139
  • sci luci1.0140-0143

06:39 Nardini program, target q5 (J0303)
  • acq 0144-00147
  • sci luci1.0148-0171
  • seeing about 0.8"-1.0""
  • 07:23 seeing 1.3"on the DIMM but the FWHM is about 1"
  • 07:36 seeing about 1.1"on the DIMM

07:44 switch to MODS

08:06 start Moretti's program, target A14
  • execMODS --mods1 A14.acq mods1.0002-0004
  • during the acquisition we put the pivot in y=724, after that the script does a blind offset
  • 08:29 execMODS --mods1 A14_sci.txt mods1.0005-0006
  • seeing about 1.5-2"on the DIMM but the FWHM is about 1.1"-1.2"
  • 09:00 seeing about 1.2"-1.4""
  • 09:02 seeing about 1.5"-1.6"
  • 09:09 seeing 1.1"-1.2" on the DIMM
  • 09:12 seeing about 1.2"-1.3"

09:16 since the seeing was very variable during the integration of the last image, we take onether mods1r.0007
  • seeing is instable
  • 09:28 seeign about 1.8"-2"

09:45 we slew to De Carli's program. * the script fort he acquisition of the mask 21 is not correct, we slew to the mask 22. mods1r.0008-0009
  • the alignment procedure failes due to the bad seeing (about 2.0") and the finding chart wrong

10:15 we slew to Antoniucci program's, target V900mon. seeing about 1.8"-1.9"
  • we change the guide star in the script and we sent it again
  • acq mods1r.0010-0014
  • sci mods1r.0015-0018, mods1b.0002-0005

11:20 slew to V899
  • seeing about 1.7"
  • acq mods1r.0019-0021
  • sci mods1r.0022-0025, mods1b.0006-0009
  • 11:29 seeing about 1.8"-2.0"
  • 11:33 seeing 1.3"-1.4"
  • 11:40 seein 1.6"
  • 11:50 seeing 1.6"-1.7"

12:05 standard for Antoniucci and Moretti, seeing bad 1.6"-2.00"
  • acq mods1r.0026-0027
  • sci mods1r.0028-0032, mods1b.0010-0012 dual grating
  • =mods1r.0031-0032 red grating

Calibration MODS

Moretti
  • flat with slit 1.0" red grating mods1r.0034-0039
  • flat with slit 5.0" red grating mods1.0040-0042
  • flat slitless red grating mods1r.0043-0047
  • lamps with slit 0.6"==mods1r.0048-0050==
  • lamps wiht slit 1.0" mods1r.0051-0053

Antoniucci
  • flat with slit 0.8" dual grating mods1r.0054-0056, mods1b.0013-0015
  • flat with slit 0.8" blue grating mods1b.0016-0018
  • flat with slit 0.8" red grating mods1r.0057-0059
  • lamps with slit 0.6" dual grating mods1r.0060-0062,mods1b.0019-0021
  • flat slitless dual grating mods1r.0063-0067, mods1b.0022-0031

Calibration LUCI

Antoniucci
  • flat Hkspec Halo2(luci1) Halo3(luci2) luci1.0177-0181,luci2.0089-0093
  • arc HKspec Neon luci1.0182-0183,luci2.0094-0095
  • arc HKspec Ar luci1.0184-0185,luci2.0096-0097
  • arc HKspec Xe luci1.0186-0187,luci2.0098-0099
  • flat zjspec Halo2 luci1.0193-0197,luci2.0105-0109
  • arc zjspec Ne luci1.0198-0199,luci2.0110-0111
  • arc zjspec Ar luci1.0200-0201luci2.0112-0113

Nardini
  • flat Ks Halo2 luci1.0203-0208
  • arc Ks Ne luci1.0209-0210
  • arc Ks Ar luci1.0211-0212
  • arc Ks Xe luci1.0213-0214
Topic revision: r12 - 19 Nov 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback